QUESTION 4
(a) Derive the emf equation for the double wound transformer. [5 marks]
(b) Describe the principle of operation of the D.C machines operating on motor mode.
[5 marks]
1
Expert's answer
2021-10-12T01:54:27-0400

"\text{Induced emf according to Lenz's law}\\
e=-N\dfrac{d\phi}{dt}\\
e-\text{Induced emf}\\
N- \text{Turns of conductors}\\
\phi-\text{magnetic flux}\\
t- time\\
\phi=\phi_m\sin{2\pi{f}t}\\
e=-N\dfrac{d(\phi_m\sin{2\pi{f}t})}{dt}\\
e=-2\pi{f}N\phi_m\cos{2\pi{f}t}\\
e_{max}=2\pi{f}N\phi_m\\
e_{rms}=\dfrac{e_{max}}{\sqrt{2}}\\
e_{rms}=\dfrac{ 2\pi{f}N\phi_m }{\sqrt{2}}\\
e_{rms}=4.44fN\phi_m\\"